OpenStudy (anonymous):

simplify e^(-3ln1/2)

OpenStudy (jamesj):

Well, what is \[ e^{\ln a} = ... \]

OpenStudy (jamesj):

\[ e^{\ln a} = a \] yes. Now what's, \( -3 \ln(1/2) \) equal to?

OpenStudy (jamesj):

\[ a \ln x = \ln x^a \]

OpenStudy (jamesj):

Hence \[ -3 \ln 1/2 = \ln (1/2)^3 \]

OpenStudy (jamesj):

**correction \[ -3 \ln 1/2 = \ln(1/2)^{-3} = \ln 2^3 \]

OpenStudy (jamesj):

Now see what to do?

OpenStudy (jamesj):

ranstucker, write your answers here, not in private message to me

OpenStudy (anonymous):

3ln2

OpenStudy (jamesj):

you could do that, but we want \[ e^{-3 \ln(1/2) } \] so it is useful to write -3 ln (1/2) as just ln(x) for some x, because then \[ e^{-3 \ln(1/2)} = e^{ln x} = x \]

OpenStudy (anonymous):

so what is the answer?

OpenStudy (jamesj):

well, if \[ -3 \ln (1/2) = \ln x \] what is x equal to?

OpenStudy (jamesj):

scroll up, we already calculated this

OpenStudy (anonymous):

I've got to get to class

OpenStudy (jamesj):

bye then